Network Fun

George Ross 9th November 1998 • Shared medium — originally a piece of coax to

which everyone was attached (later developments are basically just a piece of virtual coax).

• Listen to the wire. If it’s quiet then you can start talking. (CSMA)

• Everyone hears everything, but discards anything not meant for them.

• If someone else starts talking at the same time, then all stop, back off a (different) random amount, and then start again. (/CD)

• Like “polite conversation.”

• Collisions (when two or more stations start talking at once) are a normal part of the arbitration mechanism. Collisions are NOT bad for your network’s health.

• Data to be transmitted (IP datagrams) are handed to the ether driver in chunks of up to 1500 bytes.

• Addressing information (destination address, source address, packet type) is added.

• Whole is transmitted in DEC order as: a preamble to condition the wire and sync the receivers; the ether frame; and a CRC for error discovery.

Bit 1 if set means that the frame is intended for all recipients.

Bit 2 if set means that the address is “locally administered” rather than globally assigned.

• Original ethernet (“10base5”) used thick yellow coax. Could run up to about 500m.

• Fine for fixed installations, but not so handy where flexibility required, so “thinnet” (“cheapernet”) developed, using same technology but on thinner, lighter, cheaper coax (“10base2”). Could run 185m.

• Maximal radius determined by velocity of propagation on the wire and the requirement that a minimal-length packet correctly operate the collision mechanism.

• Absolute maximum of 1024 stations on an ethernet, imposed by the collision backoff algorithm.

• Maximum of 30 cable joins on thinnet, due to reflections from impedance discontinuities.

These limitations can be worked around in two ways:

repeaters just copy everything that comes in one port out to all their other ports. They clean up the signal, but that’s all. (They nibble off a few bits of preamble.) Repeaters propagate collisions.

bridges watch source addresses and learn which segments machines are on. They buffer incoming frames and only re-transmit them to the Right Place. Bridges terminate collision domains.

The 5–4–3 rule states that between all pairs of stations in a collision domain there can be at most:

• 5 pieces of wire or fibre

• 4 repeaters

• 3 coax segments • “10baseT” originally developed to use old phone

wire. Now uses more tightly specified wire to allow higher speeds and greater distances.

• Uses two twisted pairs (TX and RX) instead of one single wire. More twists are better.

• Differential signalling gives better noise immunity (CMR) and reduced EMI.

• A “hub” is just a multi-port repeater. A “switch” is just a multi-port bridge.

• No longer a shared physical medium, but still a shared logical medium implemented by the hubs and switches.

• Length is limited by wire losses rather than lightspeed. 10baseT will stretch to 150m on cat-5 TP.

• But wiring standards say 90m of fixed wiring and 10m for patch wires. That’s what developers are assuming when designing faster protocols (100baseT).

• There’s no 100base2 or 100base5 — coax is dead. Faster networks consist of repeaters, switches and end-stations interconnected with TP or fibre.

• 100baseT repeater rules are MUCH more restrictive than 10baseT. 5–4–3 rule becomes 3–2 rule, and the inter-repeater segment can be at most 5m.

Pro repeaters:

• Simple

• Cheap

• Diagnostic snooping easy

Con repeaters:

• Simple, brainless

• Shared bandwith on logical wire

• Security-breach snooping easy

• Repeater-count and distance limitations Pro bridges:

• Higher aggregate bandwidth

• No security-breach snooping

• Mixed speeds on same logical ether

• Terminate collision domains, easing distance problems

• More intelligence (VLANs, trunking, multi-cast filtering)

Con bridges:

• More expensive, more complicated to set up

• Diagnostic snooping less straightforward

• Senders can gang up on receivers

\overline{Why...? 9}

• The backbone is very short of bandwidth

• There are so many client wires (we keep splitting those for bandwidth) that attaching servers and routers in anything like an optimal way is tricky

• Servers are connected to their client wires using 10baseT, and so are limited by that bandwidth

• Snooping is possible on a shared medium

• Routing and IP packet filtering is being done piggy-back on our general servers, adding to their load, rather than on dedicated boxes

• We would like to extend VLANs to Appleton Tower and the rest of Informatics

• We might want protocol-based VLANs, with some routed locally and some bridged and/or routed by the EdLAN core.

• Much of our office wiring is “old” and incapable of 100baseT use